BLOCbeamlineStatus » History » Revision 2
« Previous |
| Next »
Nebrensky, Henry, 25 February 2015 17:46
Status of MICE Muon Beamline¶
- Table of contents
- Status of MICE Muon Beamline
This list is for tracking the state between BLOCs, and should be viewed before an expert is called. This is complementary to the whiteboard in the control room.
Access / Construction Work¶
- Hall is open.
- PRY construction continues - hence crane activity
Interlocks and Controls¶
- Target gate valve is open
- Beamstop is closed
Services (electricity / water / air / air-conditioning / network)¶
- No service interuptions scheduled.
- Airconditioning unit 1 is isolated, the others are operational.
- The beamline polarity is set to mixed polarity
- Control boards have been swapped between Q4 and Q5 MPS to look into ramping issue.
- The Linde Fridge is running
- The PSU is isolated from the DS, and undergoing testing to look for overtemperature trips.
- If it trips, leave as-is (do NOT reset interlocks) so that the specific channel can be identified
- Proton Absorber:
- need to further close damping valves
- Powerpack 2 has a cracked hose fitting awaiting replacement and is out of use
- No issues
Updated by Nebrensky, Henry about 8 years ago · 2 revisions